# Enhanced pedestal transport driven by edge collisionality on Alcator C-Mod and its role in regulating H-mode pedestal gradients

### v1

@@ 4 Pedestal operational space and transport drive
Particle transport and fueling depend on and affect $T_{e}$ as well as $n_{e}$. Neutral transport resulting from atomic processes has strong dependence on $T_{e}$, and intrinsic changes to plasma transport modify and are modified by kinetic profiles, including $T_{e}$ [14, 15]. Figure 3 shows characteristic $n_{e}$ and $T_{e}$ pedestal profiles at low and high $P_{\mathrm{net}}$, shown in green and purple respectively. There is a clear drop in pedestal temperature, $T_{e}^{\mathrm{ped}}$ at low $P_{\mathrm{net}}$. Figure 7 shows the phase space of these plasmas in terms of $n_{e}$ and $T_{e}$, at $\psi_{n}=0.95$, ($n_{e}^{95}$, $T_{e}^{95}$). This location is slightly inside the pedestal top where gradients are considerably smaller than in the pedestal itself, so that generally, ($n_{e}^{\mathrm{ped}},T_{e}^{\mathrm{ped}})\sim(n_{e}^{95},T_{e}^{95}$). We use this location instead of the top of the pedestal defined by the mtanh to facilitate the use of collisionality as explained below in Section 4.1. Indeed $T_{e}^{95}$ varies dramatically in this set of experiments. In fact, within error bars, these pedestals cluster at opposite corners of the operational space - at high $n_{e}^{95}$ and low $T_{e}^{95}$, and low $n_{e}^{95}$ and high $T_{e}^{95}$. For the remainder of the paper, we distinguish the former using purple and the latter using green. Purple markers are also those at low $P_{\mathrm{net}}$ and high $\Gamma_{D}$, while green ones at higher $P_{\mathrm{net}}$ also maintain lower $\Gamma_{D}$. We make a more quantitative distinction between these in the next section.
@@
Figure 7: Pedestal operational space in terms of ($n_{e}$,$T_{e}$). Purple diamonds are at high $n_{e}$ and low $T_{e}$, while green diamonds are at low $n_{e}$ and high $T_{e}$. Isobars (gray, dashed) and constant collisionality contours (black, dash-dotted) are included.
@@
Figure 8: Pedestal pressure (left) and normalized pressure gradient (right) plotted against pedestal collisionality. In both plots, there is a clear dependence of $p_{e}^{95}$ and $\alpha_{\mathrm{MHD}}$ with $\nu^{*}_{95}$, whereby higher $\nu^{*}_{95}$ inhibits large pedestal pressure or pressure gradient. Shown also are power law fits, to both high and low $\nu^{*}$ points separately. For the plot on the left, both high and low $\nu^{*}$ discharges trend in the same way, with $p_{e}^{95}\sim(\nu^{*}_{95})^{-0.5}$. For the plot on the right, $\alpha_{\mathrm{MHD}}$ decreases more strongly at low $\nu^{*}$ and more weakly at high $\nu^{*}$. Their trends for low and high $\nu^{*}$ are $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-1.1}$ and $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-0.7}$, respectively.
@@ 4.1 Pressure and collisionality
Figure 7 includes both isobars and constant collisionality contours (CCC), shown in gray, dashed and black, dash-dotted curves respectively. The curves are both calculated at $\psi_{n}=0.95$. The isobars are calculated using $p_{e}=n_{e}T_{e}$ and the CCCs are calculated using $\nu^{*}_{95}=\frac{q_{95}R_{0}\nu_{ei}}{\epsilon^{3/2}v_{\mathrm{th},e}}$, where $q_{95}$ is the safety factor at $\psi_{n}=0.95$, $R_{0}$ is the plasma major radius, $\nu_{ei}$ is the electron-ion collision frequency, $\epsilon$ is the inverse aspect ratio, and $v_{\mathrm{th},e}$ is the electron thermal velocity. $\nu_{e,i}$ depends on $Z_{\mathrm{eff}}$, the effective charge of the plasma, which is a difficult quantity to measure on C-Mod. Since these plasmas did not have extrinsic impurity seeding, $Z_{\mathrm{eff}}$ is chosen to be 1.4, a common assumption for relatively pure plasmas on C-Mod. Since all of these plasmas are at similar $B_{t}$ and $I_{P}$, there is little variation in $q_{95}$, so we calculate the CCCs from its average, $\langle q_{95}\rangle=4.7$. From Figure 7, it is apparent that these pedestals range substantially both in $p_{e}^{95}$ and $\nu^{*}_{95}$. Pedestals with low $D_{\mathrm{eff}}$ can reach up pressures up to $p_{e}^{95}=26.5$ kPa, more than twice that of pedestals with high $D_{\mathrm{eff}}$. Conversely, transport-ridden pedestals at low $p_{e}^{95}$ can reach collisionalities up to $\nu^{*}_{95}=4.5$. Despite rather continuous changes in heat and particle sources (as seen in Figure 2, there appears to be a quite clear separation in $\nu^{*}_{95}$. Green points are those at $\nu^{*}_{95}<1.5$ and the purple points are those at $\nu^{*}_{95}>1.5$.
@@
Plotting $p_{e}^{95}$ against $\nu^{*}_{95}$ as in the left panel of Figure 8 yields a relatively scatter-free correlation. Clearly, collisionality imposes rather strict regulation on the pedestal pressure, with $p_{e}^{95}\sim(\nu^{*}_{95})^{-0.5}$. In fact, it seems that collisionality not only affects the attainable $p_{e}^{\mathrm{ped}}$, but also the pressure pedestal gradient, $\nabla p^{\mathrm{ped}}$. Previous pedestal analysis on C-Mod has observed that the dimensionless pressure gradient is linked to the pedestal collisionality [18]. The right panel of Figure 8 supports this, showing the normalized pressure gradient, $\alpha_{\mathrm{MHD}}$, plotted against $\nu^{*}_{95}$, where $\alpha_{\mathrm{MHD}}=\frac{2\mu_{0}q_{95}^{2}R_{0}}{B_{T}^{2}}\nabla p_{e}$, using the maximum $\nabla p_{e}$ in the edge, not strictly co-located with the center of the $n_{e}$ pedestal. Note that most parameters in $\alpha_{\mathrm{MHD}}$ are relatively constant across these discharges, with the exception of $\nabla p_{e}$. In other words, changes in $\alpha_{\mathrm{MHD}}$ are primarily changes in $\nabla p_{e}$ in this dataset. As with $p_{e}$, $\nabla p_{e}$ is very tightly regulated by collisionality, decreasing even more strongly with $\nu^{*}_{95}$. At low $\nu^{*}_{95}$, $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-1.1}$. At high $\nu^{*}_{95}$, the dependence is slightly weaker, with $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-0.7}$. It may be that as $\nu^{*}_{95}$ increases further, $\alpha_{\mathrm{MHD}}$ reaches some minimum value required to sustain an H-mode.
@@
Weakening of $\nabla p_{e}$ can be a symptom of differences in the position of the $n_{e}$ pedestal relative to the $T_{e}$ pedestal. Since $\nabla p_{e}$ consists of both $\nabla n_{e}$ and $\nabla T_{e}$, its value is strongly related to the position of each gradient. Misalignment in the positions of maximum gradient strongly impact $\alpha_{\mathrm{MHD}}$. The relative shift of the two profiles has been previously observed and is thought to be a fueling effect, where the peak of the ionization moves radially outward, increasing density and locally cooling the plasma [18, 44, 45]. Of course, $\nabla p_{e}$ can also decrease if either of its constituent gradients decreases. In this dataset, discharges at low collisionality (green) have a relative density and temperature shift, $\Delta R_{\mathrm{n-T}}^{\mathrm{ped}}=\Delta R_{\mathrm{n^{\mathrm{ped}}}}-R_{\mathrm{T^{\mathrm{ped}}}}$ = 1.4 mm, whereas those at high collisionality (purple) have a relative shift $\Delta R_{\mathrm{n-T}}^{\mathrm{ped}}$ = 2.0 mm. On the other hand, low collisionality discharges have on average, a maximum $\nabla T_{e}$ = 123 keVm-1, whereas high collisionality discharges only have a maximum $\nabla T_{e}$ = 32 keVm-1 on average. Indeed, it appears that as $\nu^{*}_{95}$ increases, it is primarily the lower $\nabla T_{e}$ weakening that is responsible for the weakening of $\alpha_{\mathrm{MHD}}$ and the threatening of fusion performance, a product of a robust $\nabla p$.
@@ 4.2 Transport variation with pedestal parameters
It is not immediately obvious how to identify causality in these pedestals. Concurrent changes in $p_{e}$ (and $\nabla p_{e}$) as well as $\nu^{*}$ make it challenging to discern cause and effect. Regardless, it is clear from Section 3, that there are large changes to $D_{\mathrm{eff}}$, which appear to affect pedestal character substantially, most notably by limiting $n_{e}$. Observing that the discharges colored in purple are also more collisional, it is natural to question whether pedestal collisionality may be responsible for driving this increased transport. To test this hypothesis, the left panel of Figure 9 shows how $D_{\mathrm{eff}}$ at mid-pedestal, $D_{\mathrm{eff}}^{\mathrm{mid}}$, varies with $\nu^{*}_{95}$. Not surprisingly, there is an increase in $D_{\mathrm{eff}}^{\mathrm{mid}}$ as $\nu^{*}_{95}$ increases. Regardless, the increase in $D_{\mathrm{eff}}$ with $\nu^{*}_{95}$ is evident only for $\nu^{*}_{95}>2.5$. The right panel of Figure 9 shows the same variable on the ordinate, but now as a function of $\nu^{*}$ calculated at the separatrix. Plotting transport against $\nu^{*}_{sep}$ yields a smoother correlation between transport and collisionality, implying that the separatrix conditions may strongly influence pedestal gradients in these EDA H-modes. At the separatrix, $\nu^{*}$ can be almost an order of magnitude higher than at the pedestal top. Given the large gradients in the H-mode pedestal this variation is not surprising - in fact, reactors will likely require a collisional separatrix with view of divertor handling [46], even if the pedestal is collisionless. Based on projections for $n_{e}^{\mathrm{sep}}$ and $T_{e}^{\mathrm{sep}}$ found in [47, 48], $\nu^{*}_{\mathrm{sep}}$ on SPARC may range anywhere from below 1 and up to 10. Operation at even higher $n_{e}^{\mathrm{sep}}$ might bring this value closer to values observed here.
Figure 9: Inferred $D_{\mathrm{eff}}$ at mid-pedestal plotted against collisionality at $\psi_{n}=0.95$ (left) and the separatrix (right), both using the safety factor at $\psi_{n}=0.95$, $q_{95}$. Dashed lines represent power law fits to green and purple discharges independently. There is a clear separation in collisionality at the top of the pedestal. For low values of $\nu^{*}$, there is an unconvincing trend of $D_{\mathrm{eff}}^{\mathrm{mid}}\sim(\nu^{*}_{95})^{0.4}$. At high $\nu^{*}_{95}$, the trend is clearer and greater than linear - $D_{\mathrm{eff}}^{\mathrm{mid}}\sim(\nu^{*}_{95})^{1.2}$. When plotting against $\nu^{*}$ at the separatrix, as in the plot on the right, the separation on the abscissa disappears and both sets of points see even stronger trends in the mid-pedestal particle transport of $D_{\mathrm{eff}}^{\mathrm{mid}}\sim(\nu^{*}_{95})^{1.3}$.
@@ 4.3 Influence of separatrix condition on pedestal gradients
That the separatrix encodes information about gradients and confinement inside the separatrix is not a novel concept. It has been previously observed on ohmic discharges on C-Mod that conditions around the separatrix strongly influence pressure gradients and the subsequent accumulation of plasma pressure inside of the separatrix [20]. Using a similar framework of electromagnetic fluid drift turbulence, it has been shown on ASDEX Upgrade (AUG) that the ($n_{e}$, $T_{e}$) operational space at the separatrix may strongly influence the overall plasma confinement regime [49]. That work applies interchange-drift-Alfvén turbulence (interchange-DALF) theory, identifying $\alpha_{t}$, defined in Equation 7, as an important turbulence control parameter.
$\alpha_{t}=3.13\times 10^{-18}R_{\mathrm{geo}}\hat{q}_{\mathrm{cyl}}^{2}\frac{n_{e}}{T_{e}^{2}}Z_{\mathrm{eff}}$ (7)
@@
Here, $R_{\mathrm{geo}}$ is the geometric major radius, taken to be equal to the device major radius, $R_{0}$, and $\hat{q}_{\mathrm{cyl}}$ is the cylindrical safety factor, calculated according to $\hat{q}_{\mathrm{cyl}}=\frac{B_{t}}{B_{p}}\times\frac{\hat{\kappa}}{R_{\mathrm{geo}}/a_{\mathrm{geo}}}$, where $a_{\mathrm{geo}}$ is the geometric minor radius, taken equal to the device minor radius, and $\hat{\kappa}$ is the effective elongation, defined in [49].
@@
It is observed that large values of $\alpha_{t}$ can widen the pressure gradient scale length ($\lambda_{p}$) at the separatrix [50]. Note that this parameter has the same form as the previously plotted $\nu^{*}$, but includes a quadratic rather than linear dependence on the safety factor, $q$. The importance of this additional factor on $q$ was noticed earlier as well in the work on C-Mod [20]. There, it was shown that $\lambda_{p}$ organized better across different values of $q_{95}$ when including this additional factor of $q$. In the current work, the plasmas are at similar $q_{95}$, so little re-organization of data occurs when plotting against $\alpha_{t}$ instead of $\nu^{*}$. Doing so, however, points to a possible physical explanation. Figure 10 indicates that the transition between the high pressure and low pressure pedestals corresponds almost exactly to when the value of $\alpha_{t}=1$, the proposed marker for transition between drift-wave (DW) and resistive ballooning mode (RBM) turbulence in [49]. As the separatrix gets denser and colder and $\alpha_{t}$ approaches unity, RBMs may begin to drive particle transport, substantially increasing $D_{\mathrm{eff}}$.
@@
Figure 10: Inferred $D_{\mathrm{eff}}$ at mid-pedestal plotted now against $\alpha_{t}$, a collisionality-like turbulence control parameter in interchange-DALF theory. It is suggested that when $\alpha_{t}$ reaches unity, edge turbulence transitions from DALF to interchange. $\alpha_{t}=1$ corresponds fairly well to the inflection point in the trend of $D_{\mathrm{eff}}^{\mathrm{mid}}$ with $\alpha_{t}$. The trend across both regimes is similar to that with $\nu^{*}_{\mathrm{sep}}$, with $D_{\mathrm{eff}}^{\mathrm{mid}}\sim\alpha_{t}^{1.2}$.

### v2

@@ 4 Pedestal operational space and transport drive
Particle transport and fueling depend on and affect $T_{e}$ as well as $n_{e}$. Neutral transport resulting from atomic processes has a strong dependence on $T_{e}$, and intrinsic changes to plasma transport modify and are modified by kinetic profiles, including $T_{e}$ [14, 15]. Figure 3 shows characteristic $n_{e}$ and $T_{e}$ pedestal profiles at low and high $P_{\mathrm{net}}$, shown in green and purple respectively. There is a clear drop in $T_{e}^{\mathrm{ped}}$ at low $P_{\mathrm{net}}$. Figure 6 shows the phase space of these plasmas in terms of $n_{e}$ and $T_{e}$, at $\psi_{n}=0.95$, ($n_{e}^{95}$, $T_{e}^{95}$). This location is slightly inside the pedestal top where gradients are considerably smaller than in the respective pedestals themselves, so that generally, they are not very different from the value at the top of the pedestal, and so, ($n_{e}^{\mathrm{ped}},T_{e}^{\mathrm{ped}})\sim(n_{e}^{95},T_{e}^{95}$). We use this location instead of the top of the pedestal defined by the mtanh to facilitate the use of collisionality as explained below in Section 4.1. Indeed $T_{e}^{95}$ varies dramatically in this set of experiments. In fact, within error bars, these pedestals cluster at opposite corners of the operational space - at high $n_{e}^{95}$ and low $T_{e}^{95}$, and low $n_{e}^{95}$ and high $T_{e}^{95}$. For the remainder of the paper, we distinguish the former using purple and the latter using green. Purple markers are also those at low $P_{\mathrm{net}}$ and high $\Gamma_{D}$, while green ones at higher $P_{\mathrm{net}}$ also maintain lower $\Gamma_{D}$. We make a more quantitative distinction between these in the next section.
@@
Figure 6: Pedestal top operational space in terms of ($n_{e}$,$T_{e}$) at $\psi_{n}$ = 0.95. Purple diamonds are at high $n_{e}$ and low $T_{e}$, while green diamonds are at low $n_{e}$ and high $T_{e}$. Isobars (gray, dashed) and constant collisionality contours (black, dash-dotted) are included.
@@
Figure 7: Pressure at $\psi_{n}=0.95$ (top) and maximum normalized pressure gradient (bottom) plotted against collisionality at $\psi_{n}=0.95$ (left) and at the separatrix (right). In both plots at left, there is a clear dependence of $p_{e}^{95}$ and $\alpha_{\mathrm{MHD}}$ on $\nu^{*}_{95}$, whereby higher $\nu^{*}_{95}$ inhibits large pedestal pressure or pressure gradient. Plotting against $\nu^{*}_{\mathrm{sep}}$ shows similar dependence, but with a less abrupt transition in $\nu^{*}$. Shown also are power law fits, to both high and low $\nu^{*}_{95}$ points separately for the plots on the left, and across the full range of $\nu^{*}_{\mathrm{sep}}$ for those on the right. For the plot on the top left, both high and low $\nu^{*}$ discharges trend in the same way, with $p_{e}^{95}\sim(\nu^{*}_{95})^{-0.5}$. For the plot on the bottom, $\alpha_{\mathrm{MHD}}$ decreases more strongly at low $\nu^{*}$ and more weakly at high $\nu^{*}$. Their trends for low and high $\nu^{*}$ are $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-0.9}$ and $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-0.7}$, respectively. When regressing against $\nu^{*}_{\mathrm{sep}}$, as in the plots on the right, a stronger dependence is found for both $p_{e}^{95}$ and $\alpha_{\mathrm{MHD}}$, with $p_{e}^{95}\sim(\nu^{*}_{\mathrm{sep}})^{-0.8}$ and $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{\mathrm{sep}})^{-1.3}$.
@@ 4.1 Pressure and collisionality
Figure 6 includes both isobars in gray, dashed curves and constant collisionality contours (CCC) in and black, dash-dotted curves. The curves are both calculated at $\psi_{n}=0.95$. The isobars are calculated using $p_{e}=n_{e}T_{e}$ and the CCCs are calculated using $\nu^{*}_{95}=\frac{q_{95}R_{0}\nu_{ei}}{\epsilon^{3/2}v_{\mathrm{th},e}}$, where $q_{95}$ is the safety factor at $\psi_{n}=0.95$, $R_{0}$ is the plasma major radius, $\nu_{ei}$ is the electron-ion collision frequency, $\epsilon$ is the inverse aspect ratio, and $v_{\mathrm{th},e}$ is the electron thermal velocity. $\nu_{e,i}$ depends on $Z_{\mathrm{eff}}$, the effective charge of the plasma, which is a difficult quantity to measure on C-Mod. Since these plasmas did not have extrinsic impurity seeding, $Z_{\mathrm{eff}}$ is taken to be 1.4, a common assumption for relatively pure plasmas on C-Mod. Since all of these plasmas are at similar $B_{t}$ and $I_{P}$, there is little variation in $q_{95}$, so we calculate the CCCs from its average, $\langle q_{95}\rangle=4.7$. From Figure 6, it is apparent that these pedestals range substantially both in $p_{e}^{95}$ and $\nu^{*}_{95}$. Pedestals with low $D_{\mathrm{eff}}$ can reach up pressures up to $p_{e}^{95}=26.5$ kPa, more than twice that of pedestals with high $D_{\mathrm{eff}}$. Conversely, transport-ridden pedestals at low $p_{e}^{95}$ can reach collisionalities close to $\nu^{*}_{95}=10$. Despite rather continuous changes in heat and particle sources as seen in Figure 2, there appears to be a quite clear separation in $\nu^{*}_{95}$. Green points are those at $\nu^{*}_{95}<2$ and the purple points are those at $\nu^{*}_{95}>2$.
@@
Plotting $p_{e}^{95}$ against $\nu^{*}_{95}$ as in the top left panel of Figure 7 yields a relatively scatter-free correlation. Clearly, collisionality imposes rather strict regulation on the pedestal pressure, with $p_{e}^{95}\sim(\nu^{*}_{95})^{-0.5}$ across both groups of data. In fact, it seems that collisionality not only affects the attainable $p_{e}^{\mathrm{ped}}$, but also the pressure pedestal gradient, $\nabla p^{\mathrm{ped}}$. Previous pedestal analysis on C-Mod has observed that the dimensionless pressure gradient is linked to the pedestal collisionality [18]. The right panel of Figure 7 supports this, showing the normalized pressure gradient, $\alpha_{\mathrm{MHD}}$, plotted against $\nu^{*}_{95}$, where $\alpha_{\mathrm{MHD}}=\frac{2\mu_{0}q_{95}^{2}R_{0}}{B_{T}^{2}}\nabla p$. We use the maximum $\nabla p$ in the edge, not strictly co-located with the center of the $n_{e}$ pedestal. Since $T_{i}$ measurements were not available for these discharges, we take $p_{i}=p_{e}$ and hence, set $\nabla p=2\nabla p_{e}$. Furthermore, the expression used here is for a circular cross section and does not account for geometric corrections. Since the plasma shape is constant in all discharges shown here and the values of $\alpha_{\mathrm{MHD}}$ reported are not meant to make direct contact with a stability analysis but rather motivate the governing physics, we opt for this simplified version. Note that most parameters in $\alpha_{\mathrm{MHD}}$ are relatively constant across these discharges, with the exception of $\nabla p_{e}$. In other words, changes in $\alpha_{\mathrm{MHD}}$ are primarily changes in $\nabla p_{e}$ in this dataset.
As with $p_{e}$, $\nabla p_{e}$ is very tightly regulated by collisionality, decreasing even more strongly with $\nu^{*}_{95}$. At low $\nu^{*}_{95}$, $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-0.9}$. At high $\nu^{*}_{95}$, the dependence is slightly weaker, with $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{95})^{-0.7}$. It may be that as $\nu^{*}_{95}$ increases further, $\alpha_{\mathrm{MHD}}$ reaches some minimum value required to sustain an H-mode. When comparing against collisionality at the separatrix, $\nu^{*}_{\mathrm{sep}}$, both $p_{e}^{95}$ and $\alpha_{\mathrm{MHD}}$ trend more strongly across the range in $\nu^{*}_{\mathrm{sep}}$, with $p_{e}^{95}\sim(\nu^{*}_{\mathrm{sep}})^{-0.8}$ and $\alpha_{\mathrm{MHD}}\sim(\nu^{*}_{\mathrm{sep}})^{-1.3}$. They exhibit greater scatter, however, perhaps as a result of the radial separation in the parameters on the ordinate and abscissa.
@@
Weakening of $\nabla p_{e}$ can be a symptom of differences in the position of the $n_{e}$ pedestal relative to the $T_{e}$ pedestal. Since $\nabla p_{e}$ consists of both $\nabla n_{e}$ and $\nabla T_{e}$, its value is strongly related to the position of each gradient. Misalignment in the positions of maximum gradient strongly impact $\alpha_{\mathrm{MHD}}$. The relative shift of the two profiles has been previously observed and is thought to be a fueling effect, where the peak of the ionization moves radially outward, increasing density and locally cooling the plasma [18, 49, 50, 51, 52, 53]. Of course, $\nabla p_{e}$ can also decrease if either of its constituent gradients decreases. In this dataset, discharges at low $\nu^{*}$ (green) have a relative density and temperature shift, $\Delta R_{\mathrm{n-T}}^{\mathrm{ped}}=\Delta R_{\mathrm{n^{\mathrm{ped}}}}-R_{\mathrm{T^{\mathrm{ped}}}}$ = 1.4 mm, whereas those at high $\nu^{*}$ (purple) have a relative shift $\Delta R_{\mathrm{n-T}}^{\mathrm{ped}}$ = 2.0 mm. On the other hand, low $\nu^{*}$ discharges have on average, a maximum $\nabla T_{e}\approx$ 123 keVm-1, whereas high $\nu^{*}$ discharges only have a maximum $\nabla T_{e}\approx$ 32 keVm-1 on average. When multiplied by $\frac{R}{T_{e}^{\mathrm{ped}}}$ as in [54], giving an estimate of the $T_{e}$ gradient scale length, $L_{T_{e}}$ normalized by machine major radius, $R$, the discharges at low $\nu^{*}$ have an average $\frac{R}{L_{T_{e}}}\approx$ 162, while those at high $\nu^{*}$ have an average $\frac{R}{L_{T_{e}}}\approx$ 76, less than half of those at low $\nu^{*}$. While the gradients are computed at a different radial location than in [54], it is informative to note that multi-machine scalings may diverge in these very high $\nu^{*}$ pedestals. Indeed, it appears that as $\nu^{*}_{95}$ increases, it is primarily the $\nabla T_{e}$ weakening that is responsible for the weakening of $\alpha_{\mathrm{MHD}}$ and a threat to fusion performance, reliant on a strong $\nabla p$.
@@ 4.2 Transport variation with pedestal parameters
Simultaneous changes in $p_{e}$, $\nabla p_{e}$ and $\nu^{*}$ make it challenging to discern causality. Regardless, it is clear from Section 3 that there are large changes to $D_{\mathrm{eff}}$, which appear to affect pedestal character substantially, most notably by limiting $n_{e}$. Observing that the discharges colored in purple are also more collisional, it is natural to question whether pedestal collisionality may be responsible for driving this increased transport. To test this hypothesis, the left panel of Figure 8 shows how $D_{\mathrm{eff}}$ at mid-pedestal, $D_{\mathrm{eff}}^{\mathrm{mid}}$, varies with $\nu^{*}_{95}$. Not surprisingly, there is an increase in $D_{\mathrm{eff}}^{\mathrm{mid}}$ as $\nu^{*}_{95}$ increases. Regardless, the increase in $D_{\mathrm{eff}}$ with $\nu^{*}_{95}$ is evident only for $\nu^{*}_{95}>2$. The right panel of Figure 8 shows the same metric for transport on the ordinate, but now as a function of $\nu^{*}$ calculated at the separatrix. At the separatrix, $\nu^{*}$ can be almost an order of magnitude higher than at the pedestal top. Given the large gradients in the H-mode pedestal this variation is not surprising. Plotting transport against $\nu^{*}_{\mathrm{sep}}$ yields a smoother correlation between transport and collisionality, implying that the separatrix conditions may more strongly influence pedestal gradients in these EDA H-modes than those at the pedestal top.
Figure 8: Inferred $D_{\mathrm{eff}}$ at mid-pedestal plotted against collisionality at $\psi_{n}=0.95$ (left) and the separatrix (right), both using $q_{95}$. Dashed lines represent power law fits to green and purple discharges independently. There is a clear separation in collisionality at the top of the pedestal. For low values of $\nu^{*}$, there is an unconvincing trend of $D_{\mathrm{eff}}^{\mathrm{mid}}\sim(\nu^{*}_{95})^{0.3}$. At high $\nu^{*}_{95}$, the trend is clearer - $D_{\mathrm{eff}}^{\mathrm{mid}}\sim(\nu^{*}_{95})^{1.0}$. When plotting against $\nu^{*}$ at the separatrix, the separation on the abscissa disappears and both sets of points see even stronger trends in the mid-pedestal particle transport of $D_{\mathrm{eff}}^{\mathrm{mid}}\sim(\nu^{*}_{\mathrm{sep}})^{1.2}$.
@@ 4.3 Influence of separatrix condition on pedestal gradients
That the separatrix encodes information about gradients and confinement inside the separatrix is not a novel concept. It has been previously observed on ohmic discharges on C-Mod that conditions around the separatrix strongly influence pressure gradients and the subsequent accumulation of plasma pressure inside of the separatrix [20]. Using a similar framework of electromagnetic fluid drift turbulence, it has been shown on ASDEX Upgrade (AUG) that the ($n_{e}$, $T_{e}$) operational space at the separatrix may strongly influence the overall plasma confinement regime as well as proximity to density limit-induced plasma disruption [55, 56]. These works apply interchange-drift-Alfvén turbulence theory, identifying $\alpha_{t}$ as an important turbulence control parameter. The full expression for $\alpha_{t}$ can be found in Equation 10 of [57], but for these C-Mod plasmas with $Z_{\mathrm{eff}}$ = 1.4, it can be approximated by Equation 7.
$\alpha_{t}=2.98\times 10^{-18}R_{\mathrm{geo}}\hat{q}_{\mathrm{cyl}}^{2}\frac{n_{e}}{T_{e}^{2}}Z_{\mathrm{eff}}$ (7)
@@
Here, $R_{\mathrm{geo}}$ is the geometric major radius, taken to be equal to the device major radius, $R_{0}$, and $\hat{q}_{\mathrm{cyl}}$ is the cylindrical safety factor, calculated according to $\hat{q}_{\mathrm{cyl}}=\frac{B_{t}}{B_{p}}\times\frac{\hat{\kappa}}{R_{\mathrm{geo}}/a_{\mathrm{geo}}}$, where $a_{\mathrm{geo}}$ is the geometric minor radius, taken equal to the device minor radius, and $\hat{\kappa}$ is the effective elongation, defined in [55] in terms of triangularity and geometric elongation, which in this work is taken as the elongation calculated by EFIT.
@@
It is observed that large values of $\alpha_{t}$ can widen the pressure gradient scale length ($\lambda_{p}$) at the separatrix [57]. Note that this parameter has the same form as the previously plotted $\nu^{*}$, but includes a quadratic rather than linear dependence on the safety factor, $q$. The importance of this additional factor on $q$ was noticed earlier as well in the work on C-Mod [20]. There, it was shown that $\lambda_{p}$ organized better across different values of $q_{95}$ when including this additional factor of $q$. In the current work, the plasmas are at similar $q_{95}$, so little re-organization of data occurs when plotting against $\alpha_{t}$ instead of $\nu^{*}$. Doing so, however, points to a possible physical explanation. Figure 9 indicates that the transition between the high pressure and low pressure pedestals corresponds to values approaching and exceeding unity, the range in which resistive ballooning mode (RBM) turbulence is thought to exert strong influence over drift-wave (DW) in [55]. As the separatrix gets denser and colder and $\alpha_{t}$ approaches unity, RBMs may begin to drive particle transport, substantially increasing $D_{\mathrm{eff}}$.
@@
Figure 9: Inferred $D_{\mathrm{eff}}$ at mid-pedestal plotted now against $\alpha_{t}$, a collisionality-like turbulence control parameter in interchange-DALF theory. It is suggested that when $\alpha_{t}$ reaches unity, edge turbulence transitions from DALF to interchange. The trend across both regimes is the same as with $\nu^{*}_{\mathrm{sep}}$, with $D_{\mathrm{eff}}^{\mathrm{mid}}\sim\alpha_{t}^{1.2}$.